OpenStudy (anonymous):

A math teacher has a total of 88 books which consist of algebra 2 books and geometry books. She has 4 more geometry books than algebra 2 books. How many geometry books does she have? Mary wants to mail a box which has dimensions (5x-5) inches long, (3x+2) inches high and (2x) inches wide. Express the volume as a polynomial. If $2,000 is invested in an account at an interest rate of 6% compounded continuously, in about how many years will the money be worth $3200? Please help!

OpenStudy (anonymous):

let x be alg 2 books then geo books =x+4 so x+4+x=88 find x x is alg 2 so geo=x+4

OpenStudy (anonymous):

=P*(1+r/100)^x=A where p is principal r is interest x is no of years A is total amount at the end of investment